A mother is 15 years older than her daughter. if the sum of their ages is 51, how old is the mother?
Found 2 solutions by karan037, lynnlo:
Answer by karan037(2)

x be age of daughter
=
x+x+15=51
=)×=18

Answer by lynnlo(4176)

mother is 33
daughter is 18
================
total of ages=====51
15+18=33+18=51